OpenStudy (anonymous):

g(x,y)=x^4+y^3 has a critical point at (0,0). What sort of critical point it is? (hint : you can use second derivative test which is : fxx(x0,y0)fyy(x0,y0)-fxy(x0y0)^2 )

OpenStudy (dan815):

do the 2nd derivative test like they suggest, if its positive then it means its increasing

OpenStudy (dan815):

do the gradient of Z=f(x,y)

OpenStudy (dan815):

then do umm divergence of that i think

OpenStudy (dan815):

this is called the laplacian right?

OpenStudy (dan815):

or just do it like they say fxx = d^2 z/dx^2 fyy=.. fyx=...

OpenStudy (dan815):

plug inot the formula fxx*fyy-fxy^2

OpenStudy (anonymous):

the problem is when ı plug into formula it gives me 0 and the formula is useless when the outcom is zero.

OpenStudy (dan815):

thts okay

OpenStudy (anonymous):

that's not okay. formula doesnt give the right answer.

OpenStudy (dan815):

it must be a point where the inflection is also happening

OpenStudy (dan815):

so the crit point is of this form

OpenStudy (dan815):



OpenStudy (dan815):

or pointing down

OpenStudy (dan815):

since its either both positive or negative around critical points i think its called an unstable crit point

OpenStudy (anonymous):

the answer says they are saddle points?

OpenStudy (dan815):

yeah saddle points

OpenStudy (dan815):

have you seen what the graph looks like on wolfram

OpenStudy (dan815):

im not sure why they said use 2nd derivative unless there is a thrm saying all unstable crit points in 3D are saddle points

OpenStudy (dan815):

but you can see by dy dz/dx and dz/dy and findting the crit point and see how it became along the x axis and then along the y axis

OpenStudy (dan815):

and you will find along the x or y axis its increasing and decreasing on the other one

OpenStudy (dan815):